# Maximum Length RLL Sequences in de Bruijn Graph

### v1

@@ 3 Construction of Sequences of Maximum Length
The construction of the maximum length $(n,s)$-sequences is based on merging all the $(n,s)$-necklaces. There are $(n,s)$-sequences that contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. But, these sequences cannot be of maximum length and this can be proved as a consequence from the proofs of Lemmas 2.8 and 2.12.
@@
Finally, we consider a maximum length acyclic $(n,s)$-sequence. This is done by adding a short path with $(n,s)$-words, which are not contained in $(n,s)$-necklaces, to the maximum length $(n,s)$-sequence ${\cal C}$. Consider an $(n,s)$-word $(10^{s}X)$, for some $s$-ones string $X$, whose length is $n-1-s$, on an $(n,s)$-necklace with the edge $(10^{s}X)\rightarrow(0^{s}X1)$ of the associated necklace on the cycle ${\cal C}$. We start the maximum length path with the vertex $(10^{s}X)$ and continue with the edge $110^{s}X$ 1 to the vertex $(0^{s}X1)$ as on the cycle ${\cal C}$ until we reach the vertex $(10^{s}X)$ again. So far the path is the same as the cycle ${\cal C}$. Now, we continue with the $s$ edges,
@@
to obtain an acyclic $(n,s)$-sequence of maximum length, i.e., $s$ edges and $s$ vertices were added to the maximum length $(n,s)$-sequence ${\cal C}$ and the outcome is an acyclic $(n,s)$-sequence which attain the bound of Theorem 2.13. Note, that all the added vertices are on a necklace which is not an $(n,s)$-necklace, but the added vertices are represented by the $(n,s)$-words which are contained in the necklace. Note, further that the added sub-path is to the end of the cycle ${\cal C}$, while in the upper bound it was considered in its beginning, but this is equivalent.
@@ 4 Enumeration of the Number of Sequences
In this section, we will consider enumeration associated with the number of $(n,s)$-words, $(n,s)$-necklaces, and $(n,s)$-sequences related to the bounds and the construction that were presented in the previous sections. For this purpose, we define two important concepts, the rate and the redundancy of the set of sequences discussed in this paper. The rate of an $(n,s)$-sequence with maximum length $N$ will be defined as
${\cal R}(n,s)=\limsup_{n\rightarrow\infty}\frac{\log N}{n}~{}.$
@@
The algorithm which uses a de Bruijn sequence of length $2^{n}$ and each one is simulated by ’11’ and each zero by ’10’ considers a sequence of length $2^{n+1}$ with windows of length $2n$ for each required $n$-tuple. Hence, it has a high redundancy of $12n-\log 2^{n+1}=n-$ 1. Our $(n,s)$-sequences reduce this redundancy quite dramatically.
What is the exact length of a maximum length $(n,s)$-sequence? The computation of this size can be done similarly to other computations associated with constrained codes [15, 21], and asymptotic computations are done in the same way. Some of the computations can be done more accurately. For example, we will present the computations for $s=1$. Let $g_{n}$ be the number of $(n,1)$-words.
@@
$g_{n}=\frac{\left(\frac{1+\sqrt{5}}{2}\right)^{n+2}-\left(\frac{1-\sqrt{5}}{2}\right)^{n+2}}{\sqrt{5}}~{}.$
@@
The number of $(n,1)$-words in the $(n,1)$-necklaces is $g_{n}-g_{n-4}$.
@@
A necklace that is not an $(n,1)$-necklace contains $(n,1)$-words if it starts with 01 and ends with 10. In between we have an $(n-4,1)$-word and hence by definition there are $g_{n-4}$ such words. Thus, the number of $(n,1)$-words in the $(n,1)$-necklaces is $g_{n}-g_{n-4}$. ∎
Now we can combine Corollary 2.12 and the constructions in Section 3 with Lemma 4.2 to obtain the following conclusion.
@@
The length of a maximum length $(n,1)$-sequence is $g_{n}-g_{n-4}$.
@@
The length of a maximum length acyclic $(n,1)$-sequence is $g_{n}-g_{n-4}+1$.
@@
$h_{n,s}=\left\lfloor\frac{\lambda^{k+1}(\lambda-1)}{(s+2)\lambda-2(s+1)}+0.5\right\rfloor,$
@@
Lemma 4.2 can be generalized as follows.
@@
A necklace which is not an $(n,s)$-necklace contains $(n,s)$-words if such a word starts with $10^{i}$ 1 and ends with $10^{j}$, where $1\leq i,j\leq s$ and $s<i+j$. In between these two substrings we have an $(n-2-i-j,s)$-word and hence by definition, there are $h_{n-2-2s,s}$ such words if $i=j=s$. If $i=s-1$, $j=s$ or $i=s$, $j=s-1$, and $s>1$, then there are $2h_{n-1-2s,s}$ such words (for both together). If $i+j=2s-2$ and $s>2$, then $i=s-2$ or $i=s-1$ or $i=s$ and hence there are $3h_{n-2s,s}$ such words. This analysis is continued for each $s+1\leq i+j\leq 2s$, where $1\leq i,j\leq s$. Finally, if $i+j=s+1$, then $1\leq i\leq s$ and hence there are $s\cdot h_{n-s+3,s}$ such words. Thus, the number of $(n,s)$-words in the $(n,s)$-necklaces is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i,s}$. ∎
@@
The length of a maximum length $(n,s)$-sequence is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i}$.
@@
The length of a maximum length acyclic $(n,s)$-sequence is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i}+s$.
Corollary 4.10.
The rate of maximum length $(n,s)$ sequences is $\log\left|\lambda\right|$ and their redundancy is $n-\left|\lambda\right|n$, where $\lambda$ is the root of Eq. (1).
We are also interested in the number of generated cycles which is $2^{K}$, where $K$ is the number of $(n,s)$-necklaces whose weights are between $k_{1}$ and $k_{2}$. This is associated with the algorithm to merge $(n,s)$-necklaces. When $s=1$ this number is not difficult to compute.
@@
It should be noted that on one hand the set $V$ defined in Section 3 is less effective when $s=1$, but based on Lemmas 4.11 and 4.12 an efficient algorithm to construct a large set of maximum length $(n,1)$-sequences can be designed. This algorithm will be based on algorithm merge $(n,1)$-necklaces and an efficient algorithm to enumerate the associated binomial coefficient, e.g., the enumerative encoding of Cover [5].
@@ 5 Sequences with Efficient Positioning Decoding
For ordering the necklaces, it is worthwhile to know their number. The total number of necklaces of order $n$, $Z(n)$, is a well-known result [8, 13]. It involves the well-known arithmetic function, the Möbius function [17].
Theorem 5.1.
The number of necklaces of order $n$ is
$Z(n)=\frac{1}{n}\sum_{d|n}\mu(d)2^{n/d},$
where $\mu(\cdot)$ is the Möbius function.
The Lyndon words have some simple properties.
@@
The Lyndon word of a necklace starts with the longest run of zeros and ends with a one.
@@
If the Lyndon word $X_{1}$ of a necklace has a larger run of consecutive zeros than in the Lyndon word $X_{2}$ of another necklace, than $x_{1}$ has a smaller value than $X_{2}$.
@@
Let $X_{1},X_{2},X_{3},~{}\cdots,~{}X_{Z(n)}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. The concatenation of these words in this order is a de Bruijn sequence of order $n$, i.e., an Eulerian circuit in $G_{n-1}$ which is also a Hamiltonian cycle in $G_{n}$.
The next step is to show that lemma 5.4 is also true if we restrict ourselves to the Lyndon words of the $(n,s)$-necklaces.
@@
Let $X_{1},X_{2},X_{3},~{}\cdots,~{}X_{Z(n)}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. There exists an index $k$ such that $X_{1},X_{2},~{}\cdots,~{}X_{k}$ are Lyndon words which are not contained in $(n,s)$-necklaces, while $X_{k+1},X_{k+2},~{}\cdots,~{}X_{Z(n)}$ are Lyndon words which are contained in $(n,s)$-necklaces.
Proof.
Since the Lyndon words are ordered by their lexicographic value from the smallest to the largest, it follows that they are first ordered by the number of leading zeros (from the largest number to the smallest number of leading zeros) in each word. By Lemma 5.2, this number of leading zeros is also the longest run of zeros in the necklace. By Lemma 5.3, all Lyndon words that start with more than $s$ zeros are contained in necklaces that are not $(n,s)$-necklaces, while all the Lyndon words that start with at most $s$ zeros are contained in $(n,s)$-necklaces which imply the claim of the lemma. ∎
@@
Proof.
By Lemma 5.3, the Lyndon words with a smaller number of leading zeros have larger values than the ones with a larger number of leading zeros. The word $0^{s+1}1^{n-s-1}$ is a Lyndon word with the largest value among the words starting with $s+1$ zeros. The next Lyndon word in the order has $s$ leading zeros and the claim of the lemma follows. ∎
@@
Proof.
All Lyndon words after $0^{s+1}1^{n-s-1}$ start with at most $s$ zeros and end with a one and hence their necklaces cannot contain any run of more than $s$ zeros. This implies that all these words are contained in $(n,s)$-necklaces. ∎
The algorithm which generates all the Lyndon words that follow the last Lyndon word $0^{s+1}1^{n-s-1}$ which is not contained in an $(n,s)$-necklace is given as follows.
Algorithm for the lexicographic generation of Lyndon words in $(n,s)$-necklaces:
The first word is defined as $X_{0}=0^{s+1}1^{n-s-1}$ (this word is not considered for the $(n,s)$-necklaces); set $Y:=X_{0}$ and $i:=0$.
  1. (E1)
    If $Y=1^{n}$, goto (E5); otherwise set $i:=i+1$.
  2. (E2)
    Find the largest index $j$ in $Y=y_{1}y_{2}~{}\cdots~{}y_{n}$ such that $y_{j}=0$.
  3. (E3)
    Let $Z=y_{1}y_{2}~{}\cdots~{}y_{j-1}1$ and set $Y$ to be the first $n$ bits of the infinite sequence $ZZZ~{}\cdots$.
  4. (E4)
    If $j$ divides $n$, then set $X_{i}:=Z$ and goto (E1); otherwise goto (E2).
  5. (E5)
    $X_{1},X_{2},\ldots,X_{i}$ is the sequence of Lyndon words of the $(n,s)$-necklaces from the smallest to the largest.
The concatenation of $X_{1}X_{2}~{}\cdots~{}X_{i}$ is the required maximum length $(n,s)$-sequence that contains exactly all the words of the $(n,s)$-necklaces.
The correctness of the algorithm is shown in the same way as it is proved in the FKM algorithm. To find the position of a given word $v$ of length $n$ in the concatenation of the Lyndon words we have to apply the algorithm proposed by Kociumaka, Radoszewski, and Rytter [19] which finds the position of $v$ in the concatenation of all the Lyndon word of order $n$. After the position of $v$ was found we have to subtract the position of the last entry in the word $0^{s+1}1^{n-s-1}$ since our sequence starts after this last entry. It is worthwhile and more efficient in the long run to compute this position in advance and save it as it will be used in every application of the algorithm. The complexity of the decoding algorithm is the same as in [19] since the ordering of the Lyndon words is the same with the exception that we do not start from the first one but with the one after $0^{s+1}1^{n-s-1}$.

### v2

@@ 31      $i\leftarrow i+1$
Other algorithms for generating de Bruijn sequences can be also adapted to merge all the $(n,s)$-necklaces for constructing $(n,s)$-sequences. The construction of the maximum length $(n,s)$-sequences is based on merging all the $(n,s)$-necklaces. There are $(n,s)$-sequences that contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. But, these sequences cannot be of maximum length by Corollary 2.14.
@@
Finally, we consider a maximum length path in $G_{n}(s)$. This is done by adding a short path with $(n,s)$-words, which are not contained in $(n,s)$-necklaces, to the maximum length cycle ${\cal C}$ in $G_{n}(s)$. Consider an $(n,s)$-word $(10^{s}X)$, for some $s$-ones string $X$, whose length is $n-1-s$, on an $(n,s)$-necklace with the edge $(10^{s}X)\rightarrow(0^{s}X1)$ of the associated necklace on the cycle ${\cal C}$. We start the maximum length path with the vertex $(10^{s}X)$ and continue with the edge $110^{s}X$ 1 to the vertex $(0^{s}X1)$ as on the cycle ${\cal C}$ until we reach the vertex $(10^{s}X)$ again. So far the path is the same as the cycle ${\cal C}$. Now, we continue with the $s$ edges,
@@
to obtain a path of maximum length in $G_{n}(s)$, i.e., $s$ edges and $s$ vertices were added to the maximum length cycle ${\cal C}$ and the outcome is a path in $G_{n}(s)$ which attain the bound of Theorem 2.15. All the added vertices are on a necklace ${\cal X}$ which is not an $(n,s)$-necklace, but the added vertices are represented by the $(n,s)$-words which are contained in ${\cal X}$. Note, that the added sub-path is to the end of the cycle ${\cal C}$, while in the upper bound it was considered in its beginning, but this is equivalent.
Corollary 3.2.
For any $11\leq s<n-$ 1 we have that a maximum length path in $G_{n}(s)$ has length $\ell_{n,s}+s$.
@@ 4 Enumeration of the Number of Sequences
In this section, we will consider enumeration associated with the number of $(n,s)$-words, $(n,s)$-necklaces, and $(n,s)$-sequences related to the bounds and the construction that were presented in the previous sections. For this purpose, we define the two well-known important concepts, the rate and the redundancy of the set of sequences discussed in this paper. The rate of an $(n,s)$-sequence ${\mathcal{S}}$ with maximum length $N$ will be defined as
${\cal R}(n,{\mathcal{S}})=\limsup_{n\rightarrow\infty}\frac{\log N}{n}~{}.$
@@
The rate and redundancy are measures for the evaluating the amount of used bit information for the sequence (ratio for the rate and difference for the redundancy.
The scheme in [37] which was discussed in Section 1 uses a de Bruijn sequence of length $2^{n}$, where each one is simulated by ‘11’ and each zero by ‘10’. This scheme considers a sequence of length $2^{n+1}$ with windows of length $2n$ for each required $n$-tuple. Hence, it has a high redundancy of $12n-\log 2^{n+1}=n-$ 1 and its rate is 0.5 . Our $(n,s)$-sequences reduce this redundancy and increase this rate quite dramatically.
What is the exact length of a maximum length $(n,s)$-sequence? The computation of this size with a closed formula can be done similarly to other computations associated with constrained codes [22, 28], and asymptotic computations are done in the same way. Some of the computations can be done more accurately. For example, we will present the computations for $s=1$. Let $g_{n}$ be the number of $(n,1)$-words. The value of $g_{n}$ is the well-known Fibonacci number which appears extensively in the theory of constraints codes.
@@
$g_{n}=h_{n,1}=\frac{\left(\frac{1+\sqrt{5}}{2}\right)^{n+2}-\left(\frac{1-\sqrt{5}}{2}\right)^{n+2}}{\sqrt{5}}~{}.$
@@
The number of $(n,1)$-words in the $(n,1)$-necklaces is $\ell_{n,1}=g_{n}-g_{n-4}$.
@@
A necklace that is not an $(n,1)$-necklace contains $(n,1)$-words if it starts with 01 and ends with 10. In between we have an $(n-4,1)$-word and hence by definition there are $g_{n-4}$ such words. Thus, the number of $(n,1)$-words in the $(n,1)$-necklaces is $\ell_{n,1}=g_{n}-g_{n-4}$. ∎
Now we can combine Theorem 2.11 and the constructions in Section 3 with Lemma 4.2 to obtain the following conclusion.
@@
The maximum length of an $(n,1)$-sequence is $g_{n}-g_{n-4}$.
@@
The maximum length of an acyclic $(n,1)$-sequence is $g_{n}-g_{n-4}+n$, i.e., a trial of length $g_{n}-g_{n-4}+1$ in $G_{n-1}(s)$.
We can now have as a consequence that even a maximum length $(n,1)$-sequence improves the rate of the naive scheme which was used before with de Bruijn sequences for a quantum key distribution scheme.
@@
$h_{n,s}=\left\lfloor\frac{\lambda^{n+1}(\lambda-1)}{(s+2)\lambda-2(s+1)}+0.5\right\rfloor,$
@@
Some values of the maximum length $(n,s)$-sequences and the value of $h_{n,s}$ are presented in Table 1 and the rates of the sequences are presented in Table 2, where ${\cal R}(n,s)=\limsup\limits_{n\rightarrow\infty}\frac{\log N}{n}=\log\lambda$.
$s$ $n$ $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$
1 $2$ $2$ $3$ $3$ $4$ $5$ $7$ $8$ $11$ $13$ $18$ $21$ $29$ $34$ $47$ $55$ $76$ $89$ $123$ $144$ $199$ $233$
2 $4$ $4$ $7$ $7$ $11$ $13$ $21$ $24$ $39$ $44$ $71$ $81$ $131$ $149$ $241$ $274$ $443$ $504$ $815$ $927$
3 $8$ $8$ $15$ $15$ $26$ $29$ $51$ $56$ $99$ $108$ $191$ $208$ $367$ $401$ $708$ $773$ $1365$ $1490$
4 $16$ $16$ $31$ $31$ $57$ $61$ $113$ $120$ $223$ $236$ $439$ $464$ $863$ $912$ $1695$ $1793$
5 $32$ $32$ $63$ $63$ $120$ $125$ $239$ $248$ $475$ $492$ $943$ $976$ $1871$ $1936$
6 $64$ $64$ $127$ $127$ $247$ $253$ $493$ $504$ $983$ $1004$ $1959$ $2000$
7 $128$ $128$ $255$ $255$ $502$ $509$ $1003$ $1016$ $2003$ $2028$
Table 1: The maximum length of a cyclic $(n,s)$-sequence and the number of $(n,s)$-words for $1\leq n\leq 11$ and $1\leq s\leq\text{maximum}(7,n)$.
$s$ $1$ $2$ $3$ $4$ $5$ $6$
$\lambda$ $1.6180$ $1.8393$ $1.9276$ $1.9659$ $1.9836$ $1.9920$
$\log\lambda$ $0.6942$ $0.8791$ $0.9468$ $0.9752$ $0.9881$ $0.9942$
$s$ $7$ $8$ $9$ $10$ $11$ $12$
$\lambda$ $1.9960$ $1.9980$ $1.9990$ $1.9995$ $1.9998$ $1.9999$
$\log\lambda$ $0.9971$ $0.9986$ $0.9993$ $0.9996$ $0.9998$ $0.9999$
Table 2: The rate, ${\cal R}(n,s)$, of maximum length $(n,s)$-sequences for $1\leq s\leq 12$.
Lemma 4.2 can be generalized as follows for $s>1$.
@@ 7 $128$ $128$ $255$ $255$ $502$ $509$ $1003$ $1016$ $2003$ $2028$
If an $(n,s)$-word starts with a one, then after this one we can have any one of the $h_{n-1,s}$ $(n-1,s)$-words. Otherwise, an $(n,s)$-word is an $(n,s)$-necklaces has a prefix $10^{i_{1}}$ 1 and a suffix $10^{i_{2}}$, where $i_{1}\geq 1$ and $i_{1}+i_{2}=i\leq s$, i.e., $1\leq i\leq s$, and between them there could be any one of the $h_{n-i-2,s}$ $(n-i-2,s)$-words which implies the claim of the lemma. ∎
@@
The maximum length of a cyclic $(n,s)$-sequence is $\ell_{n,s}=h_{n-1,s}+\sum_{i=1}^{s}(i\cdot h_{n-i-2,s})$. The maximum length of an acyclic $(n,s)$-sequence is $\ell_{n,s}=h_{n-1,s}+\sum_{i=1}^{s}(i\cdot h_{n-i-2,s})+s+n-1$, where $s\leq n-2$.
@@
The rate of maximum length $(n,s)$-sequences is $\log\lambda$ and their redundancy is $n-\lambda n$, where $\lambda$ is the root of Eq. (1).
We might be interested in the number of $(n,s)$-necklaces of weight $k$. This can be associated with an algorithm to merge $(n,s)$-necklaces, with restricted weight (another possible constraint), especially when $s=1$. When $s=1$ this number is not difficult to compute.
@@
It should be noted that on one hand the set $V$ defined in Section 3 is less effective when $s=1$, but based on Lemmas 4.10 and 4.11 an efficient algorithm to construct a large set of maximum length $(n,1)$-sequences can be designed. This algorithm will be based on Algorithm 1 to merge $(n,1)$-necklaces and an efficient algorithm to enumerate the associated binomial coefficient, e.g., the enumerative encoding of Cover [7].
@@ 5 Sequences with Efficient Positioning Decoding
The Lyndon words have some simple properties which were used to merge all the necklaces into a de Bruijn sequence [14, 15, 16].
@@
The Lyndon word of a nonzero necklace starts with the longest run of zeros and ends with a one.
@@
If the Lyndon word $X_{1}$ of a necklace has a larger run of consecutive zeros than in the Lyndon word $X_{2}$ of another necklace, then $X_{1}$ has a smaller value than $X_{2}$.
@@
Let $X_{1},X_{2},X_{3},~{}\ldots,~{}X_{\ell}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. The concatenation of these words in this order is a de Bruijn sequence of order $n$, i.e., an Eulerian circuit in $G_{n-1}$ which is also a Hamiltonian cycle in $G_{n}$.
The next step is to show that lemma 5.3 is also true if we restrict ourselves to the Lyndon words of the $(n,s)$-necklaces. This is the simple idea which led to efficient construction of $(n,s)$-sequences from all the $(n,s)$-necklaces [17, 30, 35]. The idea is summarized in the following lemmas.
@@
Let $X_{1},X_{2},X_{3},~{}\ldots,~{}X_{\ell}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. There exists an index $k$ such that $X_{1},X_{2},~{}\ldots,~{}X_{k}$ are Lyndon words which are not contained in $(n,s)$-necklaces, while $X_{k+1},X_{k+2},~{}\ldots,~{}X_{\ell}$ are Lyndon words which are contained in $(n,s)$-necklaces.
@@
The algorithm, which generates all the Lyndon words that follow the last Lyndon word $0^{s+1}1^{n-s-1}$ which is not contained in an $(n,s)$-necklace, is presented in Algorithm 3. It follows very similar steps to the ones in Ruskey, Savage, and Wang [30].
1
Input: $n$, $s$, $X_{0}\coloneqq 0^{s+1}1^{n-s-1}$
Output: maximum length lexicographic cyclic $(n,s)$-sequence
2 Set $Y=y_{1}y_{2}~{}\cdots~{}y_{n}\leftarrow X_{0}$ and $i\leftarrow 0$
3 while $Y\neq 1^{n}$ do
4       $j\leftarrow\max\{t\in\{1,2,\dots,n\}:y_{t}=0\}$
5       $Z\leftarrow y_{1}y_{2}~{}\cdots~{}y_{j-1}1$
6       $V=v_{1}v_{2}v_{3}\dots\leftarrow ZZZ\cdots$
7       $Y\leftarrow v_{1}v_{2}~{}\cdots~{}v_{n}$
8       if $j$ divides $n$  then
9             $i\leftarrow i+1$
10             $X_{i}\leftarrow Z$
11       end if
12
13 end while
14return $X_{1}X_{2}~{}\cdots~{}X_{i}$ (each $X_{i}$ is returned when computed)
Algorithm 3 Lexicographic generation of Lyndon words in the $(n,s)$-necklaces
The output of the algorithm is a maximum length cyclic $(n,s)$-sequence. By Theorem 2.11, Lemma 5.4, and Corollaries 5.5 and 5.6, the concatenation of $X_{1}X_{2}~{}\cdots~{}X_{i}$ is the required maximum length cyclic $(n,s)$-sequence that contains exactly all the words of the $(n,s)$-necklaces. For the acyclic $(n,s)$-sequence we had to start after the word $0^{s+1}1^{n-s-1}$ and this is the word associated with the first $n$ bits of $00^{s}1^{n-s-1}$ 0 since the first $n$ bits of $X_{1}$ are zeros followed by a one. This also implies that for the acyclic $(n,s)$-sequence after $X_{i}$ we have to add $s$ zeros which add $s$ edges to the sequence (see the discussion that follows Theorem 2.11). Hence, we have the following result which is the main theorem of this section.
Theorem 5.7.
Algorithm 3 produces maximum length cyclic $(n,s)$-sequences and a maximum length acyclic $(n,s)$-sequence is generated if in the last line it will return the sequence $0^{s}1^{n-s-1}X_{1}X_{2}\dots X_{i}0^{s}$.
The correctness of the algorithm is shown in the same way as it is proved in the FKM algorithm. To find the position of a given word $v$ of length $n$ in the concatenation of the Lyndon words we have to apply the algorithm proposed by Kociumaka, Radoszewski, and Rytter [25] which finds the position of $v$ in the concatenation of all the Lyndon word of order $n$. After the position of $v$ was found we have to subtract the position of the last entry in the word $0^{s+1}1^{n-s-1}$ since our sequence starts after this last entry. It is worthwhile and more efficient in the long run to compute this position in advance and save it as it will be used in every application of the algorithm. The complexity of the decoding algorithm is the same as in [25] since the ordering of the Lyndon words is the same with the exception that we do not start from the first one but with the one after $0^{s+1}1^{n-s-1}$. Finally, the average complexity of computing the next bit is constant using the techniques and algorithms as was explained first in [31].